# Enabling Local Neural Operators to perform Equation-Free System-Level Analysis

### v1

@@ D.1 The D1Q3 Lattice Boltzmann Model
$u_{0}(x)=w\tanh{\big{(}\alpha(x-c)\big{)}}+\beta,\quad v_{0}(x)=0.12u_{0}(x),$ (44)
$w\sim\mathcal{U}(0.8,1.2),\quad\alpha\sim\mathcal{U}(0.5,1),$
$c\sim\mathcal{U}(2,18),\quad\beta\sim\mathcal{U}(-0.4,0).$
@@ D.2 Training Neural Operators: POD-RandONets
Computing the POD modes is fast, even if using around 400,000 function-data points, because it only involves the eigen decomposition of the covariance matrix $C=XX^{T}$, which is of size $82\times 82$. Let us denote this eigen decomposition by $C=V\Lambda V^{-1}$. From this decomposition, we selected the first $k=12$ components, which correspond to a POD score of $99.9\%$. Then, we use the matrix $V_{k}^{T}\in\mathbb{R}^{12\times 82}$ to project the data $X$ onto $\tilde{X}=V_{k}^{T}X$, which is subsequently used as input to the RandONet. We call such RandONets ‘POD-RandONets’.

### v2

@@ D.1 The D1Q3 Lattice Boltzmann Model
[…]
@@ D.2 Training Neural Operators: POD-RandONets
Computing the POD modes remains efficient, even with approximately 400,000 function-data points, since it only requires the eigen-decomposition of the covariance matrix $C=XX^{T}$, which has dimension $82\times 82$. We denote the eigen-decomposition of the covariance matrix by $C=V\Lambda V^{-1}$. From this, we retain the leading $k=12$ components, capturing $99.9\%$ of the total POD energy. The projection matrix $V_{k}^{T}\in\mathbb{R}^{12\times 82}$ is then used to obtain the reduced representation $\tilde{X}=V_{k}^{T}X$, which serves as input to the RandONet. We refer to such models as POD-RandONets.